Nathen is 4 years older than Henry. If the average of their ages is 28, what is the age of Nathan?
Answer by fractalier(6550)

Call their ages now, n and h. We can write
n = h + 4 and
(n + h)/2 = 28 which becomes
n + h = 56
Now substitute in the first one and get
h+4 + h = 56
2h + 4 = 56
2h = 52
h = 26 so that
n = 30